OpenStudy (anonymous):

Find the lateral area of a cone which has 32 and 14 as dimensions. Use 3.14 for pi and round the result.

Directrix (directrix):

The lateral area of a right circular cone is 1/2 times the circular base circumference times L which is the slant height.

Directrix (directrix):

LA = 1/2 * (2 pi*r) * L LA = pi*15*28 LA = ?

OpenStudy (anonymous):

I got 1, 319

Directrix (directrix):

I like exact answers where possible. I got 420*pi which approximates to 1319.4689. To the nearest integer, that is 1319 which meshes with your calculations.

OpenStudy (anonymous):

Okay cool. Thank you!

OpenStudy (anonymous):

So the answer is 1.319 rounded right?

Directrix (directrix):

Rounded to the nearer whole number, yes.

OpenStudy (anonymous):

Thank you! Good night!